Twelve different video games showing violence were observed. The duration times of violence were​ recorded, with the times​ (seconds) listed below. Assume that these sample data are used with a 0.10 significance level in a test of the claim that the population mean is greater than 80 sec. If we want to construct a confidence interval to be used for testing that​ claim, what confidence level should be used for a confidence​ interval? If the confidence interval is found to be 21.3 sec < mu<188.1 ​sec, what should we conclude about the​ claim?
77,13,586,49,0,61,208,43,183,0,2,52
The confidence level should be ??% (Do not round)
1
Expert's answer
2020-12-08T15:06:48-0500

The confidence level should be 80%

1-2a = 1-2*0.10 = 0.8


The given confidence interval contains the value of 80 sec, so there

is not sufficient evidence to support the claim that the mean is greater than 80 sec.